How many moles of aluminum chloride and sodium hydroxide would be needed to produce 7.80 moles of aluminum hydroxide
1
Expert's answer
2020-06-10T07:37:43-0400

The reaction between aluminium chloride and sodium hydroxide can be written as following:

AlCl3 + 3NaOH = Al(OH)3 + 3NaCl

As 7.80 moles of aluminium hydroxide are produced, 7.80 moles of aluminium chloride and 7.80 × 3 = 23.4 moles of sodium hydroxide are required.


Answer: 7.80 mol AlCl3 and 23.4 mol NaOH.